OpenStudy (anonymous):

What three lengths can NOT be the lengths of the sides of a triangle? a) 23 m, 17, m,14 m or b)11 m, 11 m, 12 m or c) 5 m, 7 m, 8 m or d) 21 m, 6 m, 10 m

OpenStudy (mathstudent55):

Take each set of lengths and take two lengths at a time and add them together. If in one choice, any two lengths added together is not greater than the third length, that cannot be a triangle.

OpenStudy (anonymous):

I did that I still can't find the answer

OpenStudy (mathstudent55):

For example, here are two sets of lengths: 1) 13 cm, 12 cm, 18 cm 2) 5 in., 8 in., 15 in. Let's work with set 1) First, add each set of two lengths and compare to the third length. a. 13 cm + 12 cm = 25 cm > 18 cm (ok) b. 12 cm + 18 cm = 30 cm > 13 cm (ok) c. 13 cm + 18 cm = 31 cm > 12 cm (ok) Since each two lengths added together are greater than the third side, 13 cm, 12 cm, 18 cm can be a triangle.

OpenStudy (anonymous):

they're all greater

OpenStudy (mathstudent55):

No they are not. Now let's do example 2). a. 5 in. + 8 in. = 13 in. < 15 in. (not ok) Just from that comparison, you can tell that set 2) will not make a triangle.

OpenStudy (anonymous):

I see

OpenStudy (mathstudent55):

You need to do one set at a time. Each set may need 3 additions and comparisons.

OpenStudy (mathstudent55):

Here is your first choice: a) 23 m, 17 m,14 m What is 23 m + 17 m? Compare to 14 m What is 17 m + 14 m? Compare to 23 m What is 23 m + 14 m? Compare to 17 m

OpenStudy (anonymous):

1. 40m 2. 31m 3. 37m

OpenStudy (mathstudent55):

a) 23 m, 17 m,14 m 23m + 17 m = 40 m > 14 m (ok) 17 m + 14 m = 31 m > 23 m (ok) 23 m + 14 m = 37 m > 17 m (ok) Choice a) makes a triangle. Now do the same to choices b), c), and d).

OpenStudy (anonymous):

I'm so confused

OpenStudy (mathstudent55):

You are correct. Now see that for each sum of two sides, I compared with the third side. Each sum was greater that the third side, so choice a) does make a triangle.

OpenStudy (mathstudent55):

I don't know why you are confused. You did the three sums correctly.

OpenStudy (mathstudent55):

Now for each sum, compare it tot eh side you did not add.

OpenStudy (mathstudent55):

Your three sums gave results: 1. 40 m 2. 31 m 3. 37 m Right?

OpenStudy (anonymous):

Yes

OpenStudy (mathstudent55):

Now all you need to do is compare each sum to the side that was NOT added. 1. 40 m compared to 14 m 2. 31 m compared to 31 m 3. 37 m compared to 17 m In each case, you see that the sum is greater than the third side right?

OpenStudy (anonymous):

I do

OpenStudy (anonymous):

yes all greater

OpenStudy (mathstudent55):

Great. That means these three sides CAN make a triangle. That means choice a) makes a triangle is not our answer.

OpenStudy (mathstudent55):

Now we do the same for choice b) We take each two sides at a time and add them. Then we compare the sum with the side we did not add. I'll set it up for you.

OpenStudy (anonymous):

so 22

OpenStudy (anonymous):

11 plus 11

OpenStudy (mathstudent55):

b)11 m, 11 m, 12 m Let's drop the m (for meters) just to make it quicker to write).

OpenStudy (mathstudent55):

Right. 11 + 11 compared to 12 11 + 11 = 22 > 12

OpenStudy (anonymous):

So 22 is greater than 12

OpenStudy (mathstudent55):

Since the sum is greater than the third side, it's good so far. We need to test the other possibilities.

OpenStudy (mathstudent55):

Exactly.

OpenStudy (anonymous):

so It's not going to be B correct?

OpenStudy (mathstudent55):

Now we test 11 + 12 = ? and compare with 11

OpenStudy (mathstudent55):

Correct. B is not it. Now we move on to c).

OpenStudy (mathstudent55):

c) 5 m, 7 m, 8 m 5 + 7 = ? compared to 8 7 + 8 = ? compared to 5 5 + 8 = ? compared to 7

OpenStudy (mathstudent55):

Is each sum in b) greater than the third side?

OpenStudy (anonymous):

1) 12 and 2) 15 and 3) 13

OpenStudy (anonymous):

8 is the same

OpenStudy (mathstudent55):

12 > 8 15 > 5 13 > 7 Each sum is greater than the third side. Once again, c) works.

OpenStudy (anonymous):

So now D

OpenStudy (mathstudent55):

c) works meaning it makes a triangle, so it's not our answer.

OpenStudy (anonymous):

I hate math because It takes forever to get done with just one problem

OpenStudy (mathstudent55):

Now let's look at d)

OpenStudy (anonymous):

21 m, 6m, 10m

OpenStudy (mathstudent55):

d) 21 m, 6 m, 10 m

OpenStudy (mathstudent55):

One way to shorten this type of problem is to look for a set of three sides in which one side is much larger than the other two.

OpenStudy (anonymous):

21 is much larger then 6 and 10m

OpenStudy (mathstudent55):

We just found it. Notice that 21 is much larger than 6 or 10. Wow, you're ahead of me.

OpenStudy (anonymous):

I think to much with math that's my problem

OpenStudy (mathstudent55):

6 + 10 = 16, and 16 < 21 Set d) will not make a triangle, so we found our answer.

OpenStudy (anonymous):

more then I have to

OpenStudy (anonymous):

there we go!

OpenStudy (anonymous):

thank you! you're patient, I'm terrible at math

OpenStudy (mathstudent55):

We went through a long problem, but don't think it;'s a waste of time. 1) we did it together, and now you understand it. 2) after going through 3 choice that were not it, you got to practice that method a bit. 3) we figured out a quick way to do it: look for a set of sides that has one side much larger than the other two.

OpenStudy (anonymous):

This is true, wow you're fantastic

OpenStudy (mathstudent55):

You are very welcome, and thanks. Keep practicing, studying & asking questions. That's how you learn.